Calculate the first four moments about the mean from the following frequency distribution.

          Also calculate the Moment Ratios b1 and  b2.                                               

Groups

2------4

4-------6

6------8

8-----10

10----12

12----14

14----16

16----18

18----20

Frequency

18

24

47

80

102

66

40

21

15


1
Expert's answer
2021-05-31T16:29:06-0400

The first central moment m1 is always zero.


The second central moment is:

"m_2=\frac{1}{N}\sum f_i(x_i-\overline{x})^2"

where fi is frequency,

"x_i=\{3,5,7,9,11,13,15,17,19\}"


The third central moment is

"m_3=\frac{1}{N}\sum f_i(x_i-\overline{x})^3"


The fourth central moment is

"m_3=\frac{1}{N}\sum f_i(x_i-\overline{x})^4"


Using online calculator vrcacademy.com, we get:

Number of observations:

"N=413"

Mean of X values:

"\overline{x}=10.76"


"m_2=13.76"

"m_3=3.17"

"m_4=526.43"


Moment Ratios:

skewness:

"b_1=\frac{1}{N\sigma^3}\sum (x_i-\overline{x})^3"

kurtosis

"b_2=\frac{1}{N\sigma^4}\sum (x_i-\overline{x})^4"

where "\sigma" is standard deviation.


Using online calculator www.socscistatistics.com, we get:

"b_1=0"

"b_2=-1.2"